AutoSave Off HD Unit 1 Project Fall 2020 (2) - Protected View - Excel guada gucci GG File Home Insert Draw Page Layout FormulAutoSave Off Have Unit 1 Project Fall 2020 (2) - Protected View - Excel guada gucci GG File Home Insert Draw Page Layout FormAutoSave Off Have Unit 1 Project Fall 2020 (2) - Protected View - Excel guada gucci GG File Home Insert Draw Page Layout FormAutoSave Off Have Unit 1 Project Fall 2020 (2) - Protected View - Excel guada gucci GG File Home Insert Draw Page Layout FormAutoSave Off Have Unit 1 Project Fall 2020 (2) - Protected View - Excel guada gucci GG File Home Insert Draw Page Layout Form
Answer #1
Amazing Company
Income Statement
Particulars 2019 2019 Remarks Vertical Analysis Remarks
Sales revenue 240,000 (a) 100.00% (a/a)
Less: Cost of goods sold (142,600) (b) 59.42% (b/a)
Gross profit 97,400 (c=a+b) 40.58% (c/a)
Administrative expenses
Rent expense 20,000 (d) 8.33% (d/a)
Salaries expense 48,000 (e) 20.00% (e/a)
Insurance expense 11,800 (f) 4.92% (f/a)
Legal expense 9,200 (g) 3.83% (g/a)
Depreciation expense 8,000 (h) 3.33% (h/a)
Total operating expense 97,000 (i=d+e+f+g+h) 40.42% (i/a)
Operating income 400 (j=c-i) 0.17% (j/a)
Interest expense 1,800 (k) 0.75% (k/a)
Total non operating expense 1,800 (l) 0.75% (l/a)
Profit before tax (1,400) (m=j-l) -0.58% (m/a)
Income tax expense (3,000) (n) -1.25% (n/a)
Profit after tax (4,400) (o=m+n) -1.83% (o/a)
EPS (0.18) p=o/24000 0.00% (p/a)
Common stock 240,000
Par value $10
Number of Common stock 24,000 (240000/10)
Amazing Company
Balance Statement
Particulars 2019 Remarks Vertical Analysis Remarks
Non Current asset
Property, plant and equipment 88,000 (i) Note 1 28.95% (i/viii)
Total non current assets 88,000 (ii) 28.95% (ii/viii)
Current asset
Inventory 36,000 (iii) 11.84% (iii/viii)
Accounts receivable 55,000 (iv) 18.09% (iv/viii)
Cash 120,000 (v) 39.47% (v/viii)
Prepaid insurance 5,000 (vi) 1.64% (vi/viii)
Total current assets 216,000 (vii) 71.05% (vii/viii)
Total assets 304,000 (viii) 100.00% (viii/viii)
Equity
Common Stock 240,000 (ix) 78.95% (ix/viii)
Retained earnings 24,200 (x) Note 2 7.96% (x/viii)
Total equity 264,200 (xi) 86.91% (xi/viii)
Non Current liability
Note payable 23,000 (xii) 7.57% (xii/viii)
Total non current liability 23,000 (xiii) 7.57% (xiii/viii)
Current liability
Accounts payable 12,200 (xiv) 4.01% (xiv/viii)
Salaries payable 2,600 (xv) 0.86% (xv/viii)
Sales tax payable 2,000 (xvi) 0.66% (xvi/viii)
Total current liability 16,800 (xvii) 5.53% (xvii/viii)
Total liability 39,800 (xviii) 13.09% (xviii/viii)
Total equity and liability 304,000 (xix) 100.00% (xix/viii)

Note 1

Property, plant and equipment 2019 2018
Office Equipment 53,000 47,000
Machinery & Tools 59,000 21,000
Accumulated depreciation (24,000) (16,000)
88,000 52,000

Note 2

Retained earnings 2019 2018
Retained earnings 38,600 38,600
Dividends (10,000)
Profit after tax for the year (4,400)
24,200 38,600
Amazing Company
Cash flow Statement
Particulars 2019 2019 Remarks
Cash flows from operating activities
Profit before tax (1,400)
Add: Interest expense 1,800
Add: Depreciation expense 8,000 9,800
(Increase)decrease in current assets
Inventory (20,400) (15600-36000)
Accounts receivable (20,000) (-55000+35000)
Prepaid insurance 1,000 (39,400) (-5000+6000)
Increase(decrease) in current liabilites
Accounts payable (4,600) (12200-16800)
Salaries payable (100) (2600-2700)
Sales tax payable (2,000) (6,700) (2000-4000)
Income tax paid (3,000)
Cash flows from operating activities (40,700) (a)
Cash flows from investing activities
Cash paid for purchase of office equipment (6,000)
Cash paid for purchase of tools (15,000)
(21,000) (b)
Cash flows from financing activities
Issue of common stock 80,000
Note payable paid (12,100)
Dividend paid (10,000)
Interest paid (1,800) 56,100 (c)
Net increase in cash and cash equivalents (5,600) d=(a+b+c)
Cash and cash equivalents at beginning of period 125,600 e
Cash and cash equivalents at end of period 120,000 f=d+e

Analysis of financial statement

Current ratio = current asset/current liabilities 12.86 (vii/xvii)
Working capital= current asset-current liabilities $199,200 (vii-xvii)
Cash ratio=cash and cash equivalent/current liabilities 7.14 (v/xvii)
Debt equity ratio=Debt/Shareholder's equity 0.09 (xii/xi)
Inventory turnover ratio= (cost of goods sold/average inventory)               5.53 -b/((36000+15600)/2)
Days sales in inventory=(1/Inventory turnover ratio)*365             66.04
Gross profit percentage = gross profit/sales 0.41 (c/a)
Accounts receivable turnover ratio= net credit sales/Average accounts receivable 5.33 (assuming entire sales is on credit) (a/((55000+35000)/2))
Days sales in receivable = (1/Accounts receivable turnover ratio)*365 68.44
Rate of return on total assets=profit after tax/average assets (0.02) o/((250200+304000)/2)
Assets turnover ratio= (1/Rate of return on total assets)*365 (22,987)
Rate of return on shareholder's equity=profit after tax/average equity (0.02) o/((160000+38600+264200)/2)
Earnings per share (0.18) (p)
Price/earnings ratio=market value per share/earnings per share (54.55) (10//earnings per share)
Dividend yield = annual dividend/price per share (0.04) (10000/24000)/10
Dividend payout             (2.27) (10000/o)